OpenStudy (anonymous):

Find the common ratio and the next term for the sequence shown. Express your answers as fractions or as decimals. 333 1/3, 250, 187 1/2

OpenStudy (anonymous):

333 / 250 = 1.332 If you want to check your answer follow this: 1.332 x 187.5 = approx 250. But the ratio is 1.332

OpenStudy (anonymous):

I hope this Helps! :)

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

Your Welcome ^_^ I like to help lol

OpenStudy (welshfella):

250 / 333 1/3 = 0.75

OpenStudy (welshfella):

common ratio = 0.75

OpenStudy (welshfella):

next term = 187.5 * 0.75

OpenStudy (welshfella):

the common ratio = second term / first

OpenStudy (welshfella):

if you want it as a fraction its 3/4